# Electron Magneto-Hydrodynamics in Graphene

### v1

@@ Abstract
We consider the hydrodynamic flow of an electron fluid in a channel formed in a two-dimensional electron gas (2DEG) with no-slip boundary conditions. To generate vorticity in the fluid the flow is influenced by an array of micromagnets on the top of 2DEG. We analyse the viscous boundary layer and demonstrate anti-Poiseuille behaviour in this region. Furthermore we predict a longitudinal Hall effect, where a periodic magnetic field generates a voltage modulation in the direction of transport. From the experimental point of view we propose a method for a precise measurements of the properties of different electron fluids. The results are applicable to graphene away from the charge neutrality point and to semiconductors.
@@ I Introduction
Hydrodynamic electron flow manifests itself in a range of effects that have been observed. While hydrodynamic electron transport was suggested to have been achieved already in 1995 in Ref. [5], this result was immediately challenged by Gurzhi in Ref. [7]: the conditions for hydrodynamics in the bulk material were satisfied, but they were not satisfied for hydrodynamics in a 2DEG. Reliable evidence of the hydrodynamic regime was first presented two decades later in Refs. [8, 9, 10]. Notable consequences of hydrodynamic behavior are the decrease of resistance with increasing temperature known as the Gurzhi effect [5, 11, 12, 13, 14, 15], Poiseuille flow profiles [16, 17, 18], negative non-local resistance and formation of whirlpools [19, 20, 8, 21], the formation of local viscous regions [22], Hall viscosity [14, 23], negative magnetoresistance [24, 25, 13, 26, 10, 27, 28], the violation of the Wiedemann-Franz law [^1] Although recent work in Ref. [29] suggests that the result of a well-known experimental work [30] is not hydrodynamic in nature. [30, 31, 32, 33, 34], anomalous scaling of the resistance with the channel width [31, 9], resonant photoresistance in magnetic fields [35, 36, 37, 38, 27], and a quantum-critical dynamic conductivity [39]. Furthermore, many novel phenomena have been proposed, including the elimination of the Landauer-Sharvin resistance [40], anisotropic fluids [41, 42], dynamo effect in the electron-hole plasma [43] and the prospect of hydrodynamic spin transport [44, 45, 46] inspired by experiments on liquid mercury [47, 48]. In addition to the Poiseuille-like flow expected from electron hydrodynamics, anomalous flow profiles have been predicted in the boundary layer in charge neutral graphene in a uniform perpendicular magnetic field, with this “anti-Poiseuille” flow attributed to recombination effects [49, 50, 51].
@@
In the laminar regime in a uniform magnetic field, vorticity is generated only near the boundaries of the channel in a layer of width $\sim\sqrt{l_{ee}l_{\rm mfp}}$. The flow in this “Knudsen layer” is influenced not only by the viscous fluid effects, but also by the boundary conditions that are device specific and typically not well known. Therefore it is hard to disentangle the unknown boundary effects and unknown viscous effects.
@@
In Ref. [52] we suggested a method that involved the use of micromagnets to create wiggling hydrodynamic flow and hence create vorticity throughout the bulk of the sample, not just in the boundary region. This allows for a dramatic enhancement of viscous hydrodynamic dissipation. Hydrodynamics in a finite channel is inherently a boundary problem and in Ref. [52] the problem was solved only for a straight channel in the unique case of perfect-slip boundaries. In the present work we extend the method to a straight channel with arbitrary boundary conditions, either no-slip or finite-slip boundaries. Of course the bulk solution is the same, but the structure of the Knudsen layer depends on the boundary condition and we determine the structure in the present work. Hence, the present work extends the magneto-hydrodynamic method to graphene and also to $\delta$-doped semiconductors.
The structure of this paper is as follows. In Sec. II we briefly describe the solution for the bulk flow through a sinusoidal magnetic field found in Ref. [52]. The difference from Ref. [52] is that we present numerical estimates for graphene instead of GaAs. In Sec. III we study the boundary layer in a straight channel using no-slip boundary condition and characterize the flow near the boundary. We first solve the boundary conditions using perturbation theory and then expand this idea to develop a numerically exact solution. The key results of Sec. III are the width of the boundary layer and the emergence of pseudo anti-Poiseuille flow at high magnetic field strengths. Understanding the width of the boundary layer allows a lower limit to be placed on the sample size in order to use the bulk solution for the “zigzag” flow. In Sec. VI we calculate the magnetic field of periodic ferromagnetic stripes and hence determine parameters of micromagnets necessary for the experimental set up assuming monolayer graphene with experimentally realistic density and mobility. Sec. VII presents our conclusions.
@@ II Bulk Solution for Electron Hydrodynamics in a sinusoidal magnetic field
Consider a two-dimensional channel where $x$ is the longitudinal direction, $y$ the transverse direction and $z$ is perpendicular to the plane. The electron fluid may be controlled through the Lorentz force and thus obeys the modified stationary Navier-Stokes equation at constant density [53],
$\mathbf{v}/\tau-\nu\nabla^{2}\mathbf{v}+\mathbf{v}\cdot{\nabla}\mathbf{v}$ $=-\nabla\Phi/m^{*}+q/m^{*}\mathbf{v}\times\mathbf{B},$
@@
Here, $\mathbf{v}$ is the fluid velocity field, $\nu$ is the kinematic viscosity, $\tau$ is the relaxation time related to Ohmic processes (scattering from impurities and phonons), $\Phi$ is the electro-chemical potential, $q$ is the charge of the fluid particle ($q=-e<0$ for electrons), $m^{*}$ is the effective hydrodynamic mass and $\mathbf{B}$ is the magnetic field. For a uniform field, it is known that this equation predicts the usual Hall effect, whereas recently it has been realized that non-uniform fields have remarkable consequences such as a local suppression of flow [54] and vortex layers in the bulk [52]. The electron-electron scattering length, $l_{ee}$, and the momentum diffusive scattering length, $l_{\rm mfp}$, are defined by the following relations
@@
where $g=2\pi/a$ and $a$ is the separation between adjacent micromagnets. As previously mentioned, the electron fluid is very viscous, so the “Bernoulli term”, $\mathbf{v}\cdot{\nabla}\mathbf{v}$, in the left hand side of Eq. (II) can be neglected. Hence, we obtain a linear Stokes equation that can be solved using the stream function $\psi$ defined by $v_{x}=\partial_{y}\psi$ and $v_{y}=-\partial_{x}\psi$. For an infinitely long and infinitely wide channel with fixed averaged current density $\mathbf{J}=qnv_{0}\mathbf{\hat{x}}$ the solution is
$\psi_{\infty}=v_{0}\left(y+\frac{\omega\tau}{g(1+D^{2}g^{2})}\cos(gx)\right),$ (4)
where $D=\sqrt{\nu\tau}=\frac{1}{2}\sqrt{l_{ee}l_{\rm mfp}}$ is the characteristic viscous length and $\omega=|e|B_{0}/m^{*}$ is the cyclotron frequency. The viscous and Ohmic dissipation rates follow directly from Eq. (4), were $L$ and $W$ are the length and the width of the channel, respectively,
@@
$\dot{E}_{\nu}$ $=$ $nm^{*}\nu\int d^{2}r\>\mathbf{v}\cdot\nabla^{2}\mathbf{v}$ (5)
For the bulk solution (4) this gives
$\dot{E}_{\Omega}$ $=$ $\frac{nm^{*}v_{0}^{2}}{\tau}\ \left(1+\frac{\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})^{2}}\right)LW,$
$\dot{E}_{\nu}$ $=$ $\frac{nm^{*}v_{0}^{2}}{\tau}\ \frac{g^{2}D^{2}\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})^{2}}LW.$ (6)
The coefficient $nm^{*}v_{0}^{2}LW/\tau$ can be identified as the dissipation due to the Drude resistance in a conventional conductor with diffusive transport. The ratio of viscous to Ohmic dissipation is dependent on the dimensionless parameters $\omega\tau$ and $gD$ which, for fixed $\omega\tau$, may be shown to be maximized at $gD=[1+(\omega\tau)^{2}/2]^{1/4}$.
So far we have not explicitly solved for the electro-chemical potential in the Stokes equation (II). Substituting $\mathbf{v}$ calculated from the stream function (4), it is easy to derive an expression for the electro-chemical potential, representing the external force on each fluid particle,
$\partial_{x}\Phi=-\frac{m^{*}v_{0}}{\tau}\left(1+\frac{\omega^{2}\tau^{2}\sin^{2}(gx)}{1+D^{2}g^{2}}\right),$
@@
As the electron number density is approximately constant throughout the sample, the thermodynamic pressure gradients vanish. Therefore, the electrochemical potential gradients are predominantly electric in nature, and we can write the voltage as $\varphi=-\Phi/|e|$,
$\varphi=\frac{m^{*}v_{0}}{|e|\tau}\left(x+\frac{\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})}\left(x-\frac{\sin(2gx)}{2g}\right)\right).$ (8)
From Eq. (8) it is clear that the electric field varies with $x$. As this effect arises from the presence of a periodic magnetic field and the modulation in voltage is in the same direction as transport it is appropriate to label this a “longitudinal Hall effect”.
[…]
@@
[…]
With these parameters and assuming $v_{0}=100$ m/s, corresponding to $I=0.3$ $\mu$A in a narrow $W=2$ $\mu$m channel from Ref. [8], the oscillation amplitude of the longitudinal Hall effect is approximately $0.3$ $\mu$V.

### v2

# Driving Viscous Hydrodynamics in Bulk Electron Flow in Graphene Using Micromagnets

@@ Abstract
We consider the hydrodynamic flow of an electron fluid in a channel formed in a two-dimensional electron gas (2DEG) with no-slip boundary conditions. To generate vorticity in the fluid the flow is influenced by an array of micromagnets on the top of 2DEG. We analyse the viscous boundary layer and demonstrate anti-Poiseuille behaviour in this region. Furthermore we predict a longitudinal voltage modulation, where a periodic magnetic field generates a voltage term periodic in the direction of transport. From the experimental point of view we propose a method for a boundary-independent measurement of the viscosity of different electron fluids. The results are applicable to graphene away from the charge neutrality point and to semiconductors.
@@ I Introduction
Hydrodynamic electron flow manifests itself in a range of effects that have been observed. While hydrodynamic electron transport was suggested to have been achieved already in 1995 in Ref. [5], this result was immediately challenged by Gurzhi in Ref. [7]: the conditions for hydrodynamics in the bulk material were satisfied, but they were not satisfied for hydrodynamics in a 2DEG. It was not until 2016 that reliable evidence of hydrodynamic transport was presented in Refs. [8, 9, 10]. Signatures of hydrodynamic behavior include the decrease of resistance with increasing temperature called the Gurzhi effect [5, 11, 12, 13, 14, 15], large negative magnetoresistance [^1] The viscous origin of negative magnetoresistance in high purity samples is disputed [16], therefore magnetic field dependence of viscosity is an open question. [17, 18, 13, 19, 10, 20, 21], Poiseuille flow profiles [22, 23, 24], Hall viscosity [14, 25], negative non-local resistance and formation of whirlpools [26, 27, 8, 28], the development of local viscous regions [29], the violation of the Wiedemann-Franz law [^2] Although recent work in Ref. [30] suggests that the result of a well-known experimental work [31] is not hydrodynamic in nature. [31, 32, 33, 34, 35], resonant photoresistance in magnetic fields [36, 37, 38, 39, 20], anomalous scaling of the resistance with the channel width [32, 9], and a quantum-critical dynamic conductivity [40]. Moreover, novel phenomena have been predicted, such as anisotropic fluids [41, 42], the elimination of the Landauer-Sharvin resistance [43], dynamo effect in the electron-hole plasma [44] and the possibility of hydrodynamic spin transport [45, 46, 47] inspired by experiments on liquid mercury [48, 49]. In addition to the Poiseuille-like flow expected from electron hydrodynamics, anomalous flow profiles have been predicted in the boundary layer in charge neutral graphene in a uniform perpendicular magnetic field, with this “anti-Poiseuille” flow attributed to recombination effects [50, 51, 52].
@@
In the laminar regime in a uniform magnetic field, vorticity is generated only near the boundaries of the channel in a layer of width $\sim\sqrt{l_{ee}l_{\rm mfp}}$. The flow in this boundary layer is influenced not only by the viscous fluid effects, but also by the boundary conditions that are device specific and typically not well known. Therefore it is hard to disentangle the unknown boundary effects and unknown viscous effects.
@@
In Ref. [53] we suggested a method that involved the use of micromagnets to create wiggling hydrodynamic flow and hence create vorticity throughout the bulk of the sample, not just in the boundary region. This allows for a dramatic enhancement of viscous hydrodynamic dissipation. Hydrodynamics in a finite channel is inherently a boundary problem and in Ref. [53] the problem was solved only for a straight channel in the unique case of perfect-slip boundaries. In the present work we extend the method to a straight channel with arbitrary boundary conditions, either no-slip or finite-slip boundaries. Of course the bulk solution is the same, but the structure of the viscous boundary layer depends on the boundary condition and we determine the structure in the present work. Hence, the present work extends the magneto-hydrodynamic method to graphene and also to $\delta$-doped semiconductors. The results of this work allow for the viscosity of the electron fluid to be extracted from a measurement of longitudinal resistivity of a finite straight channel in such devices.
The structure of this paper is as follows. In Sec. II we describe the solution for the bulk flow through a sinusoidal magnetic field found in Ref. [53]. The difference from Ref. [53] is that we present numerical estimates for graphene instead of GaAs. In Sec. III we study the boundary layer in a straight channel using no-slip boundary condition and characterize the flow near the boundary. We first solve the boundary conditions using perturbation theory and then expand this idea to develop a numerically exact solution. The key results of Sec. III are the width of the boundary layer and the emergence of pseudo anti-Poiseuille flow at high magnetic field strengths. Understanding the width of the boundary layer allows a lower limit to be placed on the sample size in order to use the bulk solution for the “zigzag” flow. In Sec. IV we present numerical solution of the model developed in the previous section and show plots of dissipation and velocity profile. In Sec. V we complement our previous numerical solution with a conventional finite element model and qualtitatively reproduce the flow profile. In Sec. VI we calculate the magnetic field of periodic ferromagnetic stripes and hence determine parameters of micromagnets necessary for the experimental set up assuming monolayer graphene with experimentally realistic density and mobility. Sec. VII presents our conclusions.
@@ II Bulk Solution for Electron Hydrodynamics in a sinusoidal magnetic field
We consider a channel that is sufficiently wide such that boundary layers are much thinner than the channel width, allowing for the bulk flow to be analysed indendently of boundary conditions. In this regime the fluid motion of the bulk may be analysed in the limit of an infinitely wide channel.
Consider a two-dimensional channel where $x$ is the longitudinal direction, $y$ the transverse direction and $z$ is perpendicular to the plane. The electron fluid may be controlled through the Lorentz force and thus obeys the stationary Navier-Stokes equation at constant density [54],
$\frac{\mathbf{v}}{\tau}-\nu\nabla^{2}\mathbf{v}+\mathbf{v}\cdot{\nabla}\mathbf{v}$ $=-\frac{\nabla\Phi}{m^{*}}+\frac{q}{m^{*}}\mathbf{v}\times\mathbf{B},$
@@
Here, $\mathbf{v}$ is the fluid velocity field, $\tau$ is the relaxation time related to momentum non-conserving scattering from impurities and phonons, $\nu$ is the kinematic viscosity, $\Phi$ is the electro-chemical potential, $m^{*}$ is the effective hydrodynamic mass, $q$ is the fluid particle charge and $\mathbf{B}$ is the magnetic field. For a uniform field, it is known that this equation predicts the usual Hall effect, whereas recently it has been realized that non-uniform fields have remarkable consequences such as a local suppression of flow [55] and vortex layers in the bulk [53]. The electron-electron scattering length, $l_{ee}$, and the momentum diffusive scattering length, $l_{\rm mfp}$, are defined by the following relations
@@
The Navier-Stokes Eq.(II) is valid in the limit when $l_{ee}$ is much smaller than all geometric scales in the problem. There are two kinds of subleading corrections to the Navier-Stokes Eq. [10, 56]. The correction of the first kind, the dependence of viscosity on magnetic field, is parametrically proportional to $(l_{ee}/r_{c})^{2}$, where $r_{c}=mv_{F}/(eB)$ is the cyclotron radius and the correction of the second kind, the Hall viscosity, is parametrically proportional to $(l_{ee}/a)^{2}$, where $a$ is the characteristic geometric scale. We start from the correction of the the second kind which leads to the following modification of the Navier-Stokes equation
$\frac{\mathbf{v}}{\tau}-\nu\nabla^{2}\mathbf{v}+\mathbf{v}\cdot{\nabla}\mathbf{v}=\left(1+\frac{l_{ee}^{2}}{2}\nabla^{2}\right)\mathbf{F}$ (3)
$\mathbf{F}=-\frac{\nabla\Phi}{m^{*}}+\frac{q}{m^{*}}\mathbf{v}\times\mathbf{B}$
The term $(l_{ee}^{2}/2)\nabla^{2}$ acting on the Lorentz force, $(q/m^{*})\mathbf{v}\times\mathbf{B}$, is literally the “Hall viscosity” term from Ref.[10]. The term $(l_{ee}^{2}/2)\nabla^{2}$ acting on the gradient of electrochemical potential, $-\nabla\Phi/m^{*}$, can be easily derived using the method used in Ref.[10]. The electrochemical potential is a sum of electrostatic part and the pressure part, $\Phi=\Phi_{E}+\Phi_{pr}$. Strictly speaking the $(l_{ee}^{2}/2)\nabla^{2}$ correction acts only on the electrostatic part. Since $\Phi_{E}\gg\Phi_{pr}$ we disregard this very small difference. Note that the term $(l_{ee}^{2}/2)\nabla^{2}\mathbf{F}$ is time even unlike the dissipative viscosity term $\nu\nabla^{2}\mathbf{v}$ which is time odd. As this modified equation is the Navier-Stokes plus the leading correction towards the kinetic equation, we call Eq.(3) the NS+ equation.
@@
[…]
$\psi_{\infty}$ $=$ $v_{0}\left(y+\frac{C\omega\tau}{g(1+D^{2}g^{2})}\cos(gx)\right),$
$C$ $=$ $2\displaystyle 1-g^{2}l_{ee}^{2}/$ 2 (5)
where $D=\sqrt{\nu\tau}=\frac{1}{2}\sqrt{l_{ee}l_{\rm mfp}}$ is the characteristic viscous length and $\omega=|e|B_{0}/m^{*}$ is the cyclotron frequency. The viscous and Ohmic dissipation rates follow directly from Eq. (II), were $L$ and $W$ are the length and the width of the channel, respectively,
@@
$\dot{E}_{\nu}$ $=$ $-nm^{*}\nu\int d^{2}r\>\mathbf{v}\cdot\nabla^{2}\mathbf{v}$ (6)
Note that generally for the incompressible Navier-Stokes equation the expression for viscous dissipation may be explicitly split up into a bulk stress term and a boundary stress term[54], however for a periodic driving force the boundary term vanishes over an integer number of periods and vanishes when the flow at the inlet and outlet becomes uniform. The boundary stress term is also negligible when the channel length $L\to\infty$. We consider the long channel limit here, but it is also simple to ensure the other two cases are met via design of the device. For the bulk solution (II) Eqs.(II) give
$\dot{E}_{\Omega}$ $=$ $\frac{nm^{*}v_{0}^{2}}{\tau}\ \left(1+\frac{C^{2}\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})^{2}}\right)LW,$
$\dot{E}_{\nu}$ $=$ $\frac{nm^{*}v_{0}^{2}}{\tau}\ \frac{C^{2}g^{2}D^{2}\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})^{2}}LW.$ (7)
These Eqs. differ from a simlar solution in Ref. [53] by account of the leading gradient expansion correction, $C=1\to C=1-g^{2}l_{ee}^{2}/2$. Eqs.(II) determine longitudinal resitance of a long and sufficiently wide channel. If the channel has about 50 micromagnets along the length then the dissipation in the contacts is negligible compared to the bulk dissipation (not more than 1-2%) and such a channel is said to be sufficiently long. The sufficient width is a more delicate issue and this question is addressed in Sec. III where we discuss the solution for the boundary layer. Hence we propose to measure the longitudinal resistance of a sufficiently long and sufficiently wide conducting channel. The diffusive mean free time $\tau$ can be can be determined from the measurement at zero magnetic field, Further measurements of the dependence of the resistance on the amplitude of the magnetic field allows the extraction of $l_{ee}$ and hence allows one to determine the viscosity $\nu$. Explicitly, this is achieved by taking the ratio of the total resistivity and the Drude resistivity,
$\frac{\rho_{Tot}}{\rho_{D}}=\frac{\dot{E}_{\Omega}+\dot{E}_{\nu}}{\dot{E}_{\Omega}(\omega=0)}=1+\frac{C^{2}\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})}.$ (8)
So far we have not explicitly solved for the electro-chemical potential in the Stokes equation (II). Substituting $\mathbf{v}$ calculated from the stream function (II), it is easy to derive an expression for the electro-chemical potential, representing the external force on each fluid particle,
$\partial_{x}\Phi=-\frac{m^{*}v_{0}}{\tau}\left(1+\frac{C^{2}\omega^{2}\tau^{2}\sin^{2}(gx)}{1+D^{2}g^{2}}\right),$
@@
As the electron number density is approximately constant throughout the sample [^3] See discussion leading to and following Eq. (12), the thermodynamic pressure gradients vanish. Therefore, the electrochemical potential gradients are predominantly electric in nature, and we can write the voltage as $\varphi=-\Phi/|e|$,
$\varphi=\frac{m^{*}v_{0}}{|e|\tau}\left(x+\frac{C^{2}\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})}\left(x-\frac{\sin(2gx)}{2g}\right)\right).$ (10)
From Eq. (10) it is clear that the electric field varies with $x$. As this effect arises from the presence of a periodic magnetic field and the modulation in voltage is in the same direction as transport it is appropriate to label this a “magnetic field driven longitudinal voltage modulation”.
Electron transport in high-mobility graphene may be hydrodynamic in the temperature range $150\>\mathrm{K}\leq T<300\>\mathrm{K}$ [3]. To be specific we will present numerical estimates for a monolayer graphene device with mobility $\mu\approx 2\times 10^{5}$ cm2///Vs and density $n\approx 10^{12}$ cm-2 at temperature $T=220$ K with parameters similar to the device used in Ref. [8],
@@
For a magnetic field amplitude $B_{0}=75$ mT this gives $\omega\tau=1.2$. Here we keep in mind that the hydrodynamic mass for monolayer graphene is $m^{*}=p_{F}/v_{F}$, with Fermi momentum $p_{F}=\hbar\sqrt{n\pi}$, accounting for spin and valley degeneracy in graphene. With these parameters the characteristic viscous length is $D\approx 0.34\>\mu\mathrm{m}$.
Ohmic dissipation (II) is dependent on the dimensionless parameters $\omega\tau$ and $gD$ which, for fixed $\omega\tau$ , may be shown to be maximized at $gD=[1+(C\omega\tau)^{2}/2]^{1/4}$. We take $g=2.6$ $\mu$m-1 which is slightly smaller than the optimal value and which corresponds to lattice spacing $a=2.5$ $\mu$m. At a larger value of g (smaller a) the gradient non-Stokes correction $l_{ee}^{2}\nabla^{2}/2\to l_{ee}^{2}g^{2}/2$ is too large. At $a=2.5$ $\mu$m the correction is $l_{ee}^{2}g^{2}/2=0.19$.
With these parameters and assuming $v_{0}=100$ m/s, corresponding to $I=0.15$ $\mu$A per $\mu$m of channel width from Ref. [8], the oscillation amplitude of the longitudinal voltage modulation is approximately $0.3$ $\mu$V. Of course the amplitude can be increased by increasing the current.
So far in the analysis we have assumed the electron fluid is incompressible as the number density is made approximately constant via a gate voltage. To check the validity of this assumption let us compare the gate voltage $\varphi_{g}$ that determines the electron density with the longitudinal voltage drop $\varphi$.
$\varphi_{g}\approx\frac{ned}{\epsilon},\ \ \ \ \varphi\sim\frac{m^{*}v_{0}}{|e|\tau}L$ (12)
Here $d$ is the depth from the gate to the 2DEG, $\epsilon$ is the dielectric constant and $L$ is the length of the device between Ohmic contacts, Taking parameters from (II) and assuming a single sided hBN substrate with dielectric constant $\epsilon_{hBN}=4.4$[8] such that $\epsilon=2.7$ and assuming $d=50$ nm, the gate voltage may be estimated as $\varphi_{g}\approx 3.4$ V. For channel length $L=100$ $\mu$m the voltage drop along the device is $\varphi\approx 0.6$ mV. Hence $\delta n/n\sim\varphi/\varphi_{g}\sim 2\times 10^{-4}$. The incompressible approximation is valid.